• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:00:23 on March17,2005

    Entry number 141750

    keyword=Day Shift Summary

    08:00 Begin shift during run 2741 - slanted carbon.
    08:00 MCC reports problems with RF cavity. No beam.
    08:10 Beam is back.
    08:40 Begin beam studies - move hall to Controlled Access
    10:25 Going to Restricted Access. Need to remove the manlift from the Hall
    11:25 MCC says they will need/want to put beam in the Hall in one hour
    ~13:00 back to beam permit
    14:15 MCC has stuck valve problem (switchyard beamline vacuum valve won't open). 14:30 beam delivery will be delayed
    14:38 MCC says 1209Beam is coming
    15:10 did harp scan (see halog 141747). Requested 40 micro amps and started production on slanted Carbon


    Run summary
    2741 07:48-08:41 Production on Slanted Carbon
    2742 10:42- Cosmics
    2743 12:28- Cosmics with some pulsed beam

    2744 15:15- Production on Slanted Carbon